Circular No. 7206

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999co IN ANONYMOUS GALAXY S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-

ian Center for Astrophysics, report that a spectrum of SN 1999co (cf. IAUC 7205), obtained by M. Calkins with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph) on June 20.5 UT, shows it to be a type-II supernova at an early epoch. The spectrum exhibits a blue continuum with P-Cyg Balmer lines and He I absorption. Narrow emission lines of Hα and [O III] yield a recession velocity for the host galaxy of 9260 km/s. The photospheric expansion velocity of the supernova is 9000 km/s, measured from the absorption minima of Hβ and Hγ. A. V. Filippenko, University of California at Berkeley, reports that a CCD spectrum of SN 1999co, obtained by D. Stern on June 20 UT with the Shane 3-m reflector at Lick Observatory, exhibits prominent Hα emission, indicating a type-II event; the P-Cyg profiles are well developed.

SUPERNOVA 1999cp IN NGC 5468 Jha et al. write: “A spectrum of SN 1999cp, obtained by M. Calkins

on June 19.2 UT with the F. L. Whipple Obs. 1.5-m telescope (+ FAST spectrograph), reveals it to be a type-Ia supernova well before maximum light. Narrow Hα emission from the host galaxy yields a recession velocity of 2870 ± 20 km/s for NGC 5468. The supernova photospheric expansion velocities of Ca II H+K and Si II (rest 635.5 nm) are 19 700 km/s and 14 100 km/s, respectively, implying that SN 1999cp is ∼ 10 days before maximum light.” Filippenko adds (from a spectrum taken as above by Stern): “SN 1999cp is of type Ia, at least a week before maximum brightness, with strong Si II absorption near 610 nm. Continued monitoring of SN 1999cp through the coming bright run is encouraged, as it may reach mag 14.”

V382 VELORUM K. Mukai and J. Swank, Goddard Space Flight Center, write: “The

Rossi X-ray Timing Explorer performed a public TOO observation of V382 Vel during May 26.22–26.26 UT, ∼ 3 days after optical maximum. In the band 2.5–10 keV, a nominal excess over the estimated background of 0.58 ± 0.11 counts/s was obtained in four Proportional Counter Units (PCUs); however, given the fluctuations in the cosmic x-ray background, this should not be considered a detection. As an approximate upper limit, 0.8 counts/s in four PCUs would correspond to 2.5 × 10^{−12} erg cm^{−2} s^{−1} in the band 2.5–10 keV, significantly below the BeppoSAX and ASCA detections.”

1999 June 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green